a.      Let A and B and C be sets, prove that A∩(BUC) = (A∩B)U( A∩C). 


1
Expert's answer
2021-03-10T11:16:10-0500

Let "x \in A \cap \left( {B \cup C} \right)" , then "x \in A" and "x \in B \cup C" . Then "x \in A" and "x \in B" or "x \in C". Then "x \in A" and "x \in B" or "x \in A" and "x \in C", but then "x \in \left( {A \cap B} \right) \cup \left( {A \cap C} \right)" , from where "A \cap \left( {B \cup C} \right) \subset \left( {A \cap B} \right) \cup \left( {A \cap C} \right)"

Let "x \in \left( {A \cap B} \right) \cup \left( {A \cap C} \right)" . Then "x \in A" and "x \in B" or "x \in A" and "x \in C". Then "x \in A" and "x \in B" or "x \in C", then "x \in A" and "x \in B \cup C" , but then "x \in A \cap \left( {B \cup C} \right)", from where "A \cap \left( {B \cup C} \right) \supset \left( {A \cap B} \right) \cup \left( {A \cap C} \right)" .

Since "A \cap \left( {B \cup C} \right) \subset \left( {A \cap B} \right) \cup \left( {A \cap C} \right)" and "A \cap \left( {B \cup C} \right) \supset \left( {A \cap B} \right) \cup \left( {A \cap C} \right)" then "A \cap \left( {B \cup C} \right) = \left( {A \cap B} \right) \cup \left( {A \cap C} \right)" .

The statement is proven